### v3

@@ 6.1. Good-for-MDP automata
Let ${\mathcal{M}}=\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\rangle$ be an MDP and ${\mathcal{A}}=\langle\Sigma,Q,q_{0},\delta,\alpha\rangle$ be a nondeterministic automaton. When the player takes actions in ${\mathcal{M}}$ the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated. Assume that, simultaneously, the player also generates a run of ${\mathcal{A}}$. This process can be modeled by an MDP that is a product of ${\mathcal{M}}$ with ${\mathcal{A}}$, and in each step the player takes an action of ${\mathcal{M}}$ and chooses a successor state in ${\mathcal{A}}$ that extends the current run on the new letter drawn by ${\mathcal{M}}$. Formally, we define the MDP ×${\mathcal{M}}\times{\mathcal{A}}=\langle S^{\prime},\langle s_{0},\bot\rangle,(A_{\langle s,q\rangle})_{\langle s,q\rangle\in S^{\prime}},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{A}}},\Sigma,\gamma\rangle$ that is defined as follows. The set of states of ×${\mathcal{M}}\times{\mathcal{A}}$ is $S^{\prime}=(S\times Q)\cup\{\langle s_{0},\bot\rangle\}$, for $\bot\notin Q$. For all $\langle s,q\rangle\in S\times Q$, the set of actions $A_{\langle s,q\rangle}$ available from $\langle s,q\rangle$ consists of all pairs $\langle a,q^{\prime}\rangle\in A_{s}\times\delta(q,\tau(s))$. In addition, the actions available from $\langle s_{0},\bot\rangle$ are $A_{\langle s_{0},\bot\rangle}=A_{s_{0}}\times\{q_{0}\}$. Then, for all $\langle s,\ast\rangle\in S^{\prime}$ and $\langle a,q^{\prime}\rangle\in A_{\langle s,\ast\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{A}}}(\langle s,\ast\rangle,\langle a,q^{\prime}\rangle,\langle s^{\prime},q^{\prime}\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$. Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{A}})\cup\Omega({\mathcal{M}}\times{\mathcal{A}})$, $r=\langle\langle s_{0},\bot\rangle,\langle s_{1},q_{0}\rangle,\langle s_{2},q_{1}\rangle,\ldots\rangle$, we define $\pi_{\mathcal{A}}(r)=\langle q_{0},q_{1},q_{2},\ldots\rangle$, and $\pi_{\mathcal{M}}(r)=\langle s_{0},s_{1},s_{2},\ldots\rangle$ to be the ${\mathcal{A}}$-run and ${\mathcal{M}}$-run that correspond to $r$ respectively. Conversely, for $r_{\mathcal{M}}=\langle s_{0},s_{1},s_{2},\ldots\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{A}}=\langle q_{0},q_{1},q_{2},\ldots\rangle\in(Q_{\mathcal{A}})^{*}\cup(Q_{\mathcal{A}})^{\omega}$ of the same length, let $r_{\mathcal{M}}\oplus r_{\mathcal{A}}=\langle\langle s_{0},\bot\rangle,\langle s_{1},q_{0}\rangle,\langle s_{2},q_{1}\rangle,\ldots\rangle$. Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{A}}\in\Pi({\mathcal{M}}\times{\mathcal{A}})\cup\Omega({\mathcal{M}}\times{\mathcal{A}})$, we have $\pi_{\mathcal{A}}(r_{\mathcal{M}}\oplus r_{\mathcal{A}})=r_{\mathcal{A}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{A}})=r_{\mathcal{M}}$.
@@
The semantic satisfaction probability of a strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ with respect to ${\mathcal{A}}$, denoted $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$, is defined as:
$\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{\mathcal{M}}^{\mu}(r\in\Omega({\mathcal{M}},\mu)|\tau(r)\in L({\mathcal{A}})).$
@@
That is, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)$ is the probability that when the player plays with $\mu$, the word $\tau(r)$ generated by ${\mathcal{M}}$ is in $L({\mathcal{A}})$. Then, the semantic satisfaction probability of an MDP ${\mathcal{M}}$ with respect to ${\mathcal{A}}$, denoted $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$, is defined by
$\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}})}\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu).$
That is, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ is the least upper bound on how good the player can do when playing with ${\mathcal{M}}$ and ${\mathcal{A}}$. It is well known that for every nondeterministic automaton ${\mathcal{A}}$ with an $\omega$-regular acceptance condition, the suprimum is attained by a pure finite-memory strategy[11].
@@
Similarly, the syntactic satisfaction probability of $\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ and ${\mathcal{M}}$, denoted $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu)$ and $\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}$ respectively, are defined by
$\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{{\mathcal{M}}\times{\mathcal{A}}}^{\mu}(r\in\Omega({\mathcal{M}}\times{\mathcal{A}},\mu)|\pi_{\mathcal{A}}(r)\mbox{ is accepting}),$
@@
$\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})}\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu).$
As with the semantic case, the value $\mathrm{PSyn}^{\mathcal{A}}_{{\mathcal{M}}}$ is attained. Moreover, it is guaranteed to be attained by a pure memoryless strategy [11].
In general, $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$, but equality is not guaranteed, as the optimal resolution of nondeterministic choices may require access to future events.
An automaton ${\mathcal{A}}$ is Good-For-MDPs (GFM, for short) if for every MDP ${\mathcal{M}}$, it holds that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$.
@@ 6.2. The nondeterminism hierarchy and GFMness
In this section we locate GFMness in the nondeterminism hierarchy. For the semantic hierarchy, it is proved in [13] that GFM-NBWs are as expressive as NBWs. More specifically, every NBW can be translated into an equivalent GFM-NBW. Unlike the case of co-Büchi automata, where nondeterministic and deterministic automata have the same expressive power, making the expressiveness equivalence of GFM-NCWs and NCW straightforward, NBWs are more expressive than DBWs, making the result about GFM-NBWs interesting. Another aspect that has been studied is deciding GFMness [30]. In  [30], the authors prove an exponential-time upper bound as well as a polynomial-space lower bound for the problem of deciding whether an automaton is GFM-NBW.
We study the syntactic hierarchy, we first point out that every HD automaton is GFM. Then, by relating GFMness and almost-DBPness, we show that SD-NBWs are GFM, yet SD-NCWs need not be GFM. Thus, with respect to the syntactic hierarchy, GFM-NBWs are between SD and nondeterministic Büchi automata, while GFM-NCWs are incomparable to SD-NCWs.
@@
Consider an HD automaton ${\mathcal{A}}$. Let $f$ be an HD strategy for ${\mathcal{A}}$. Then, every strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ can be augmented with $f$ to obtain a strategy $\mu\times f\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSem}^{\mathcal{A}}_{{\mathcal{M}}}(\mu)=\mathrm{PSem}^{\mathcal{A}}_{{\mathcal{M}}\times{\mathcal{A}}}(\mu\times f)=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu\times f)$. Thus, $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}\leq\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$, and so ${\mathcal{A}}$ is GFM.
@@
Consider an $\omega$-regular automaton ${\mathcal{A}}$ over an alphabet $\Sigma$, and consider the MDP ${\mathcal{M}}=\langle\Sigma,\sigma_{0},(A_{\sigma})_{\sigma\in\Sigma},\mathrm{Pr}_{{\mathcal{M}}},\mathsf{id}\rangle$, where $\sigma_{0}\in\Sigma$ is an arbitrary initial state, $A_{\sigma}=\{\ast\}$ is a dummy set of actions for each letter $\sigma\in\Sigma$, and $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\sigma^{\prime})=|\Sigma|^{-1}$, for all $\sigma,\sigma^{\prime}\in\Sigma$. By [11], there is a pure memoryless strategy $\mu^{*}:S\times Q_{\mathcal{A}}\cup\{\langle s_{0},\bot\rangle\}\rightarrow A\times Q_{\mathcal{A}}$ that attains $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}$. Observe that since $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\cdot)=(\sigma^{\prime}\mapsto|\Sigma|^{-1})$ is the uniform distribution on $\Sigma$ for all $\sigma\in\Sigma$, then ${\mathcal{M}}$ is essentially an MC that generates a random word $w=\sigma_{1},\sigma_{2},\sigma_{3},\ldots\in\Sigma^{\omega}$, such that all its letters are independent and distributed uniformly. Thus, the membership $w\in L({\mathcal{A}})$, where $w$ is the word generated by ${\mathcal{C}}({\mathcal{M}},\mu^{\prime})$ (or by ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{A}},\mu^{\prime})$) for $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}})$ (respectively $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$), is independent of the strategy $\mu^{\prime}$. In particular $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$ equals the probability that a random word $w$ is in $L({\mathcal{A}})$. Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\langle\ast,q_{0}\rangle$; i.e., there is no choice at the initial state of ×${\mathcal{M}}\times{\mathcal{A}}$. Also, as $A_{\sigma}=\{\ast\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{A}}\rightarrow Q_{\mathcal{A}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{A}}(q,\sigma)$ for all $\langle\sigma,q\rangle\in\Sigma\times Q_{\mathcal{A}}$. Thus, $\mu^{*}$ induces a pruning ${\mathcal{A}}_{\mu^{*}}$ of ${\mathcal{A}}$, and $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{A}}_{\mu^{*}}$. Now, if ${\mathcal{A}}$ is GFM, then $\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}$. Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{A}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{A}})$. That is, ${\mathcal{A}}_{\mu^{*}}$ is a deterministic pruning of ${\mathcal{A}}$ that loses only a set of words of measure zero. Thus, ${\mathcal{A}}$ is almost-DBP, and we are done.
@@
Recall that in Theorem 11, we argued that there is an SD-NCW that is not almost-DBP. By Theorem 15, we can conclude that the SD-NCW described there, is also not GFM. Thus, while HDness implies GFMness, for co-Büchi automata, SDness need not imply GFMness. In the rest of this section we prove that for Büchi automata, SDness does imply GFMness, thus all SD-NBWs are GFM. In particular, Theorem 15 suggests an alternative proof of Theorem 10, which states that all SD-NBWs are almost-DBP.
@@
If ${\mathcal{A}}$ is an SD-NBW and $\mu\in{\mathcal{F}}({\mathcal{M}})$ is a finite memory strategy, then there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{A}})$ such that $\mathrm{PSem}^{\mathcal{A}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{A}}_{\mathcal{M}}(\mu^{\prime})$.

### v4

@@ 6.1. Good-for-MDP automata
Let ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{\mathcal{M}},\Sigma,\tau\right\rangle$ be an MDP and ${\mathcal{N}}=\left\langle\Sigma,Q,q_{0},\delta,\alpha\right\rangle$ be a nondeterministic automaton. When the player takes actions in ${\mathcal{M}}$ the infinite word $\tau(r)\in\Sigma^{\omega}$ is generated. Assume that, simultaneously, the player also generates a run of ${\mathcal{N}}$. This process can be modeled by an MDP that is a product of ${\mathcal{M}}$ with ${\mathcal{N}}$, and in each step the player takes an action of ${\mathcal{M}}$ and chooses a successor state in ${\mathcal{N}}$ that extends the current run on the new letter drawn by ${\mathcal{M}}$. Formally, we define the MDP ×${\mathcal{M}}\times{\mathcal{N}}=\left\langle S^{\prime},\left\langle s_{0},\bot\right\rangle,A^{\prime},\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}},\Sigma,\gamma\right\rangle$ that is defined as follows. The set of states of ×${\mathcal{M}}\times{\mathcal{N}}$ is $S^{\prime}=(S\times Q)\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}$, for $\bot\notin Q$. The set of actions $A^{\prime}=A\left\langle s_{0},\bot\right\rangle\cup\bigcup_{\left\langle s,q\right\rangle\in S\times Q}(A_{\left\langle s,q\right\rangle})$ is defined as follows. For all $\left\langle s,q\right\rangle\in S\times Q$, the set of actions $A_{\left\langle s,q\right\rangle}$ available from $\left\langle s,q\right\rangle$ consists of all pairs $\left\langle a,q^{\prime}\right\rangle\in A_{s}\times\delta(q,\tau(s))$. In addition, the actions available from $\left\langle s_{0},\bot\right\rangle$ are $A_{\left\langle s_{0},\bot\right\rangle}=A_{s_{0}}\times\left\{q_{0}\right\}$. Then, for all $\left\langle s,\ast\right\rangle\in S^{\prime}$ and $\left\langle a,q^{\prime}\right\rangle\in A_{\left\langle s,\ast\right\rangle}$, we have $\mathrm{Pr}_{{\mathcal{M}}\times{\mathcal{N}}}(\left\langle s,\ast\right\rangle,\left\langle a,q^{\prime}\right\rangle,\left\langle s^{\prime},q^{\prime}\right\rangle)=\mathrm{Pr}_{{\mathcal{M}}}(s,a,s^{\prime})$ and $\gamma(s,\ast)=\tau(s)$. Given a finite or infinite run $r\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, $r=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$, we define $\pi_{\mathcal{N}}(r)=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle$, and $\pi_{\mathcal{M}}(r)=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle$ to be the ${\mathcal{N}}$-run and ${\mathcal{M}}$-run that correspond to $r$ respectively. Conversely, for $r_{\mathcal{M}}=\left\langle s_{0},s_{1},s_{2},\ldots\right\rangle\in S^{*}\cup S^{\omega}$ and $r_{\mathcal{N}}=\left\langle q_{0},q_{1},q_{2},\ldots\right\rangle\in(Q_{\mathcal{N}})^{*}\cup(Q_{\mathcal{N}})^{\omega}$ of appropriate lengths, let $r_{\mathcal{M}}\oplus r_{\mathcal{N}}=\left\langle\left\langle s_{0},\bot\right\rangle,\left\langle s_{1},q_{0}\right\rangle,\left\langle s_{2},q_{1}\right\rangle,\ldots\right\rangle$. Notice that for all $r_{\mathcal{M}}\oplus r_{\mathcal{N}}\in\Pi({\mathcal{M}}\times{\mathcal{N}})\cup\Omega({\mathcal{M}}\times{\mathcal{N}})$, we have $\pi_{\mathcal{N}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{N}}$ and $\pi_{\mathcal{M}}(r_{\mathcal{M}}\oplus r_{\mathcal{N}})=r_{\mathcal{M}}$.
@@
We can now formally define the notion an automaton being good-for-MDP. We use the formalism of [HPS+20]. The semantic satisfaction probability of a strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ with respect to ${\mathcal{N}}$, denoted $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$, is defined as:
$\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{\mathcal{M}}^{\mu}(r\in\Omega({\mathcal{M}},\mu):\tau(r)\in L({\mathcal{N}})).$
@@
That is, $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)$ is the probability that when the player plays with $\mu$, the word $\tau(r)$ generated by ${\mathcal{M}}$ is in $L({\mathcal{N}})$. Then, the semantic satisfaction probability of an MDP ${\mathcal{M}}$ with respect to ${\mathcal{N}}$, denoted $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$, is defined by
$\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}})}\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu).$
That is, $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ is the least upper bound on how good the player can do with respect to the objective of generating a word in $L({\mathcal{N}})$.
@@
Similarly, the syntactic satisfaction probability of $\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ and ${\mathcal{M}}$, denoted $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu)$ and $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ respectively, are defined by
$\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu)={\mathrm{Pr}}_{{\mathcal{M}}\times{\mathcal{N}}}^{\mu}(r\in\Omega({\mathcal{M}}\times{\mathcal{N}},\mu):\pi_{\mathcal{N}}(r)\mbox{ is accepting}),$
@@
$\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\sup_{\mu\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})}\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu).$
Maximizing the syntactic satisfaction can be expressed as a simple stochastic parity game that is played on top of the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, where the winning objective is inherited from ${\mathcal{N}}$. In  [CH03], the authors prove that simple stochastic parity games admit optimal pure memoryless strategies. Thus, the supremum in the definition of $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure memoryless strategy. Note that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$ is evaluated with respect to the MDP ×${\mathcal{M}}\times{\mathcal{N}}$, and thus by pure memoryless we mean that there is a strategy $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$ such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}$.
Similarly, the semantic satisfaction can be expressed as a simple stochastic parity game that augments ×${\mathcal{M}}\times{\mathcal{N}}$ with a deterministic automaton ${\mathcal{D}}$ for $L({\mathcal{N}})$. Thus, the supremum in the definition of $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}$ is attained by a pure finite-memory strategy. Indeed, the considerations are as in the syntactic case, except that here, the current state of the deterministic automaton ${\mathcal{D}}$ is used as a memory for the player.
In general, $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$. In the syntactic case, the player not only aims to generate a word in $L({\mathcal{N}})$, but also to simultaneously generate an accepting run of ${\mathcal{N}}$ on it. Therefore, any strategy for the syntactic case can be used for the semantic case by ignoring the component of ${\mathcal{N}}$, resulting in a potentially higher semantic satisfaction. Equality, however, is not guaranteed. Indeed, the same way not all automata are history deterministic, it may not be possible to resolve the nondeterminism in ${\mathcal{N}}$ in a way that maximizes the probability of generating a word in $L({\mathcal{N}})$ (thus reaching a maximal semantic satisfaction) while generating rejecting runs on words in $L({\mathcal{N}})$ with zero probability (and thus reaching an equal syntactic satisfaction). When the equality holds for all MDP ${\mathcal{M}}$, we say that ${\mathcal{N}}$ is good-for-MDP (GFM for short). Note that when ${\mathcal{N}}$ is GFM, it holds that for all MDP ${\mathcal{M}}=\left\langle S,s_{0},(A_{s})_{s\in S},\mathrm{Pr}_{{\mathcal{M}}},\Sigma,\tau\right\rangle$, there exists a pure-memroyless strategy for ×${\mathcal{M}}\times{\mathcal{N}}$, $\mu^{*}_{\mathcal{N}}:S\times Q\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q$, such that $\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSyn}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$ and $\mu^{*}_{\mathcal{N}}$ generates a rejecting run on words in $L({\mathcal{N}})$ with zero probability. Note that $\mu^{*}_{\mathcal{N}}$ can be viewed as a pure finite-memory strategy for ${\mathcal{M}}$ for which $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}(\mu^{*}_{\mathcal{N}})$.
@@ 6.2. The nondeterminism hierarchy and GFMness
In this section we locate GFMness in the nondeterminism hierarchy. For the semantic hierarchy, it is proved in [HPS+20] that GFM-NBWs are as expressive as NBWs. More specifically, every NBW can be translated into an equivalent GFM-NBW. Unlike the case of co-Büchi automata, where nondeterministic and deterministic automata have the same expressive power, making the expressiveness equivalence of GFM-NCWs and NCW straightforward, NBWs are more expressive than DBWs, making the result about GFM-NBWs interesting. Another aspect that has been studied is deciding GFMness [STZ22]. In  [STZ22], the authors prove an exponential-time upper bound as well as a polynomial-space lower bound for the problem of deciding whether an automaton is GFM-NBW.
We study the syntactic hierarchy, we first point out that every HD automaton is GFM [KMBK14]. Then, by relating GFMness and almost-DBPness, we show that SD-NBWs are GFM, yet SD-NCWs need not be GFM. Thus, with respect to the syntactic hierarchy, GFM-NBWs are between SD and nondeterministic Büchi automata, while GFM-NCWs are incomparable to SD-NCWs.
@@
The theorem is proved in [KMBK14]. We give a short proof for completeness. Consider an HD automaton ${\mathcal{N}}$. Let $f$ be an HD strategy for ${\mathcal{N}}$. Then, every strategy $\mu\in{\mathcal{F}}({\mathcal{M}})$ can be augmented with $f$ to obtain a strategy $\mu\times f\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}}(\mu)=\mathrm{PSem}^{\mathcal{N}}_{{\mathcal{M}}\times{\mathcal{N}}}(\mu\times f)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu\times f)$. Thus, $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}\leq\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$, and so ${\mathcal{N}}$ is GFM.
@@
Consider an $\omega$-regular automaton ${\mathcal{N}}$ over an alphabet $\Sigma$, and consider the MDP ${\mathcal{M}}=\left\langle\Sigma,\sigma_{0},(A_{\sigma})_{\sigma\in\Sigma},\mathrm{Pr}_{{\mathcal{M}}},\mathsf{id}\right\rangle$, where $\sigma_{0}\in\Sigma$ is an arbitrary initial state, $A_{\sigma}=\left\{\ast\right\}$ is a dummy set of actions for each letter $\sigma\in\Sigma$, and $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\sigma^{\prime})=|\Sigma|^{-1}$, for all $\sigma,\sigma^{\prime}\in\Sigma$. By [CJH03], there is a pure memoryless strategy $\mu^{*}:S\times Q_{\mathcal{N}}\cup\left\{\left\langle s_{0},\bot\right\rangle\right\}\rightarrow A\times Q_{\mathcal{N}}$ that attains $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}$. Observe that since $\mathrm{Pr}_{{\mathcal{M}}}(\sigma,\ast,\cdot)=(\sigma^{\prime}\mapsto|\Sigma|^{-1})$ is the uniform distribution on $\Sigma$ for all $\sigma\in\Sigma$, then ${\mathcal{M}}$ is essentially an MC that generates a random word $w=\sigma_{1},\sigma_{2},\sigma_{3},\ldots\in\Sigma^{\omega}$, such that all its letters are independent and distributed uniformly. Thus, the membership $w\in L({\mathcal{N}})$, where $w$ is the word generated by ${\mathcal{C}}({\mathcal{M}},\mu^{\prime})$ (or by ${\mathcal{C}}({\mathcal{M}}\times{\mathcal{N}},\mu^{\prime})$) for $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}})$ (respectively $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$), is independent of the strategy $\mu^{\prime}$. In particular $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$ equals the probability that a random word $w$ is in $L({\mathcal{N}})$.
Note that it must be that $\mu^{*}(\sigma_{0},\bot)=\left\langle\ast,q_{0}\right\rangle$; i.e., there is no choice at the initial state of ×${\mathcal{M}}\times{\mathcal{N}}$. Also, as $A_{\sigma}=\left\{\ast\right\}$, for all $\sigma\in\Sigma$, we can think of $\mu^{*}$ as a function $\mu^{*}:\Sigma\times Q_{\mathcal{N}}\rightarrow Q_{\mathcal{N}}$, with $\mu^{*}(\sigma,q)\in\delta_{\mathcal{N}}(q,\sigma)$ for all $\left\langle\sigma,q\right\rangle\in\Sigma\times Q_{\mathcal{N}}$. Thus, $\mu^{*}$ induces a pruning ${\mathcal{N}}_{\mu^{*}}$ of ${\mathcal{N}}$, and $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})$ equals to the probability that a random word is accepted by ${\mathcal{N}}_{\mu^{*}}$.
Now, since $\mu^{*}$ is an optimal strategy for ×${\mathcal{M}}\times{\mathcal{N}}$ and ${\mathcal{N}}$ is GFM, it follows that $\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{*})=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}=\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}$. Hence, the probability that a random word $w\in\Sigma^{\omega}$ is accepted by ${\mathcal{N}}_{\mu^{*}}$ equals the probability that $w\in L({\mathcal{N}})$. That is, ${\mathcal{N}}_{\mu^{*}}$ is a deterministic pruning of ${\mathcal{N}}$ that loses only a set of words of measure zero and hence ${\mathcal{N}}$ is almost-DBP.
@@
Recall that in Theorem 11, we argued that there is an SD-NCW that is not almost-DBP. By Theorem 15, we can conclude that the SD-NCW described there, is also not GFM. Thus, while HDness implies GFMness for all accaptence conditions [KMBK14], for co-Büchi automata, SDness need not imply GFMness. In the rest of this section we prove that for Büchi automata, SDness does imply GFMness, thus all SD-NBWs are GFM. In particular, Theorem 15 suggests an alternative proof of Theorem 10, which states that all SD-NBWs are almost-DBP.
@@
If ${\mathcal{N}}$ is an SD-NBW and $\mu\in{\mathcal{F}}({\mathcal{M}})$ is a finite memory strategy, then there exists $\mu^{\prime}\in{\mathcal{F}}({\mathcal{M}}\times{\mathcal{N}})$ such that $\mathrm{PSem}^{\mathcal{N}}_{\mathcal{M}}(\mu)=\mathrm{PSyn}^{\mathcal{N}}_{\mathcal{M}}(\mu^{\prime})$.